OpenStudy (anonymous):

Really need help understanding this limit/derivative question... The limit Lim x approachs 5pi CosX +1 / x-5pi represents the derivative of some function f(x) at some number a. Find f and a . I don't even understand the wording of this question...

zepdrix (zepdrix):

Remember the limit definition of a derivative?\[\large f'(x)=\lim_{h \rightarrow 0}\frac{f(x-h)-f(x)}{h}\]Well there is also another definition that we see less often, of this form,\[\large f'(x)=\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}\]This is the one that we want to analyze.

zepdrix (zepdrix):

If we compare this to the limit we were given, we can see that is looks like our A value will be 5pi yes? \[\large \lim_{x \rightarrow 5\pi}\frac{\cos x +1}{x-5\pi}\]

zepdrix (zepdrix):

The top is still a little tricky though, we have to sort it out.

OpenStudy (anonymous):

wait the second one just looks like the mean value theorem? it oddly looks like f(a)-b/a-b = f'(c)

OpenStudy (anonymous):

oh nevermind me theres a limit sorry im a bit tired

zepdrix (zepdrix):

heh, yah it looks similar :)

zepdrix (zepdrix):

\[\large \cos(5\pi)=?\]

OpenStudy (anonymous):

So yeah, but we're looking for the original function is that it or looking for the derivative? and cos5pi would be... .96 ? But I think i did it in degree so it might be wrong.

zepdrix (zepdrix):

That's one of your special angles that you're going to want to remember. It will produce the same value as Pi. 5pi is Pi with an extra spin around the circle. -1 yes?

zepdrix (zepdrix):

With 2 extra spins* my bad.

OpenStudy (anonymous):

I'll ask you one thing before we continue, when taking pi in a derivative, do we ALWAYS use it in radians? I mean sometimes it works when I don't put it in my calculator as a radian

zepdrix (zepdrix):

If you're dealing with Pi, then yes you need to be in radians :o You could convert to degrees if radians are confusing you though.

OpenStudy (anonymous):

i mean pi i mean cos and sin... wow I' m sounding stupid

zepdrix (zepdrix):

5pi is the same as 180 degrees.

OpenStudy (anonymous):

so in the question am I using L'hospital rule or finding the limit? that's where I'm lost. the question is really confusing . And alright thanks for explaining the pi. :)

zepdrix (zepdrix):

No L'Hop. We're relating a weird looking limit back to the Limit Definition of a Derivative. We need to match up the pieces so we can see what the original function was. So far we've established that our A value is 5pi. If you're unsure about that, compare the form of our limit with the Definition,\[\large \large f'(x)=\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}\qquad\qquad \rightarrow \qquad \qquad \lim_{x \rightarrow 5\pi}\frac{\cos x +1}{x-5\pi}\]

zepdrix (zepdrix):

See the a?

OpenStudy (anonymous):

Yeah it's represented by 5pi?

OpenStudy (anonymous):

so i have to equate both limit? and finding my h?

zepdrix (zepdrix):

no h, we're using the second definition that i posted, not the one involving h.

zepdrix (zepdrix):

It's another form of the limit definition. It comes up less often.

OpenStudy (anonymous):

My bad I understood we had to manipulate it back to the other form

OpenStudy (anonymous):

Then when its ask to find F, am I suppose to find cosx?

OpenStudy (anonymous):

sorry, the wording is really throwing me off

zepdrix (zepdrix):

If we can show that the limit matches the DEFINITION, then we can show what our F is. So we've established that 5pi matches the A we're looking for. We've also shown that cos(5pi)=-1. If we can somehow find a -1 in the top of that fraction, we can make it look like the Definition.

OpenStudy (anonymous):

Oh wow. I understand. This was a very basic question but I have never seen that definition in my entire class. Thanks a lot! :)

zepdrix (zepdrix):

\[\large \frac{\cos x+1}{x-5\pi} \qquad =\qquad \frac{\cos x-(-1)}{x-5\pi} \qquad = \qquad \frac{\cos x-(\cos 5\pi)}{x-5\pi}\]

zepdrix (zepdrix):

Make sense? :) k cool!

OpenStudy (anonymous):

Thanks a lot you're patient!